OpenStudy (anonymous):

What does it mean to say a=b is a rational number in least terms? Compute 1-1/(2-1/7) as a rational number in least terms.

OpenStudy (anonymous):

This is a really good question Gummage I hope it gets an answer soon.

OpenStudy (anonymous):

thanks friend

OpenStudy (anonymous):

it means that it can't be simplified any further eg 1/3 whereas 2/6 is not laest terms. 1 - 1 /(13/7) = 1 - 7/13 = 6/13

OpenStudy (anonymous):

i assumed you meant a/b

OpenStudy (anonymous):

yes, thanks jimmy

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

is that literally it? are you sure? the question is 10 marks so seems a bit simple for the whole question